OpenStudy (anonymous):

describe all the seventh roots of 16 in the complex plane.

OpenStudy (phi):

magnitude 4, equi-spaced in theta in increments of 2pi/7

OpenStudy (anonymous):

Not 4...

OpenStudy (anonymous):

just it should be magnitude 16 right?

OpenStudy (phi):

oops 16^(1/7)

OpenStudy (mr.math):

\[16^{\frac{1}{7}}=\large{16^{\frac{1}{7}}e^ {\frac{2k\pi i}{7}}}\], where k=0,1,2,..6.

OpenStudy (anonymous):

^ yup